OpenStudy (rmrjr22):

suppose the velocity of a moving particle is v(t)=t^2-11t+24 ft/sec find. a) net distance traveled in the first 10 seconds b) total distance traveled in 10 seconds

OpenStudy (mani_jha):

For calculating total distance dx/dt=t^2-11t+24 dx=(t^2-11t+24)dt(Equation 1) Integrate from O to 10. For calculatng net distance, observe that \[t ^{2}-11t+24=(t-3)(t-8)\] Speed becomes 0 at t=3 and t=8. after t=3, the function becomes negative(put t=4 or 5) The body moves in opposite direction, until speed becomes zero again at t=8. To find the net distance calculate integral of equation 1 from 0 to 3 first. Then calculate integral from 3 to 8. Subtract the two. Now, after t=8, function becomes positive. Find the integral from 8 to 10, and add this to the previous result. I think its correct. Pls tell me if you have problem understanding

OpenStudy (rmrjr22):

Thanks! I believe i Got it.